The mass of a Grade A Gala apple is Normally distributed with mean 205 grams and standard deviation 5 grams. M1A1 packages of apples are intended to have no more than 2000 grams, but can exceed that mass 4% of the time. What is the maximum number of these Gala apples that should be placed into one M1A1 package, so that the mass limit is exceeded less than 4% of the time. Justify your answer by giving the probability of your value of n and also for n+1.
Answer #1

Let X be the mass of Grade A Gala Apple,

μ=205

5

Thus, X ~ N(205, 52)

Let Y be the sum of a sample of n Grade A Gala Apple then,

Y =N(n 205, n*52)  ( As per Reproductive property of the normal distribution)

So, P(Y >= 2000) < 0.04

Z = (Y-n*mu)/(sqrt{n}*sigma)

For p = 0.04 on std. normal curve

Z = 1.75

To satisfy the above condition,

Z <= 1.75

(Y-n*mu)/(sqrt{n}*sigma) <= 1.75

Soving the above equation,

n < 9.62

That is the maximum number of these Gala apples that should be placed into one M1A1 package is 9.

For n = 9,

z = (2000-205*9)/9 * 5

z = 10.33

Probability of P(Z>=10.33) < .00001

while for n = 10

z = (2000-205*10)/V10 *5

z = -3.16

Probability of P(Z>=-3.16) = 0.9992

Hence, the maximum number of these Gala apples that should be placed into one M1A1 package so that the mass limit is exceeded less than 4% of the time is 9.